# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ C.1.2 Bounded Deviation of the Generated Distribution
We start the proof by clarifying the notation for the Euler sampling approximation process. We denote its discretization timesteps by $10=t_{0}<t_{1}<\ldots<t_{K}=$ 1, with $\Delta t_{k}=t_{k}-t_{k-1}$. It is initiated at the same limit distribution as the groundtruth CTMC, $p_{\epsilon}$, and the bound to be proven will quantify the deviation that the approximated procedure incurs in comparison to the groundtruth CTMC. To accomplish so, we define $\mathcal{R}^{\theta,E}_{k}$ as the Markov kernel that corresponds to apply Euler sampling with the approximated rate matrix $R^{\theta}_{t}$, moving from $t_{k-1}$ to $t_{k}$. Therefore, $\mathcal{R}^{\theta,E}=\mathcal{R}^{\theta,E}_{1}\,\mathcal{R}^{\theta,E}_{2}\,\ldots\,\mathcal{R}^{\theta,E}_{K}$ and $p(y^{1:D}_{K})=p_{\epsilon}\mathcal{R}^{\theta,E}$.
We first apply the same decomposition to the left-hand side of Theorem 3, as Campbell et al. (2022), Theorem 1:
$\|p(y^{1:D}_{K})-p_{\text{data}}\|_{\text{TV}}$ $=\|p_{\epsilon}\mathcal{R}^{\theta,E}-p_{\text{data}}\|_{\text{TV}}$
$\leq\|p_{\epsilon}\mathcal{R}^{\theta,E}-p_{\epsilon}\mathbb{P}_{1|0}\|_{\text{TV}}+\|p_{\epsilon}-\underbrace{p(z^{1:D}_{0})}_{=p_{\epsilon}}\|_{\text{TV}}$ (21)
$\leq\|p_{\epsilon}\mathcal{R}^{\theta,E}-p_{\epsilon}\mathbb{P}_{1|0}\|_{\text{TV}}$
@@
where, in Eq. 21, $\mathbb{P}_{1|0}$ denotes the path measure of the exact groundtruth CTMC and the difference between limit distributions (second term from Eq. 21) is zero since in flow matching the convergence to the limit distribution via linear interpolation is not asymptotic (as in diffusion models) but actually attained at $t=0$. In Eq. 22, we introduce the stepwise path measure, i.e., $\mathcal{P}_{k}=\mathbb{P}_{t_{k}|t_{k-1}}$, such that $\mathbb{P}_{T|0}=\mathcal{P}_{1}\mathcal{P}_{2}\ldots\mathcal{P}_{K}$. Therefore, finding the intended upper bound amounts to establish bounds on the total variation distance for each interval $[t_{k-1},t_{k}]$.
@@
$\|\nu\mathcal{R}^{\theta,E}_{k}-\nu\mathcal{P}_{k},\|_{\text{TV}}$ $\leq\|\nu\mathcal{R}^{\theta,E}_{k}-\nu\mathcal{R}^{\theta}_{k}+\nu\mathcal{R}^{\theta}_{k}-\nu\mathcal{P}_{k}\|_{\text{TV}}$
$\leq\|\nu\mathcal{P}_{k}-\nu\mathcal{R}^{\theta}_{k}\|_{\text{TV}}+\|\nu\mathcal{R}^{\theta}_{k}-\nu\mathcal{R}^{\theta,E}_{k}\|_{\text{TV}},$ (23)
@@
For the first term, we use Proposition 5 from Campbell et al. (2022) to relate the total variation distance imposed by the Markov kernels with the difference between the corresponding rate matrices:
$\|\nu\mathcal{P}_{k}-\nu\mathcal{R}^{\theta}_{k}\|_{\text{TV}}$$\leq$ $\int_{t_{k-1}}^{t_{k}}\sup_{z^{1:D}_{t}\in\mathcal{Z}^{D}}\left\{\sum_{z^{1:D}_{t+\mathrm{d}t}\neq z^{1:D}_{t}}\left|R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})-R_{t_{k-1}}^{\theta}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})\right|\right\}\mathrm{d}t$
@@
where, in Eq. 24, we use the Mean Value Theorem, with $t_{c}\in(t_{k-1},t_{k})$; in Eq. 25, we use the fact that there are $ZD$ values of $z^{1:D}_{t+\mathrm{d}t}$ that differ at most in only one coordinate from $z^{1:D}_{t}$; in Eq. 26, we use the result from Proposition 7 to upper bound the time derivative of the multivariate unconditional rate matrix; and finally, in Eq. 27, we define $B_{k}=\sup_{t\in(t_{k-1},t_{k})}B_{t}$ and $\Delta t_{k}=t_{k}-t_{k-1}$.
@@
$U^{z^{1:D}_{t}\to z^{1:D}_{t+\mathrm{d}t}}_{k}=\sqrt{C_{0}+C_{1}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta,d}_{1|t}(z^{d}_{1}|z^{1:D}_{t})\right]},$
@@
It remains to bound the second term from Eq. 23. We start by analyzing the Markov kernel corresponding to a Markov chain with constant rate matrix $R^{\theta}_{t_{k-1}}$ between $t_{k-1}$ and $t_{k}$. In that case, from Eq. 19 we obtain:
$\mathcal{R}^{\theta}_{k}(t_{k-1},t_{k})$ $=e^{R^{\theta}_{t_{k-1}}\Delta t_{k}}$
@@
$\mathcal{R}^{\theta,E}_{k}(t_{k-1},t_{k})$ $=\tilde{p}_{t_{k}|t_{k-1}}(z^{1:D}_{t_{k}}|z^{1:D}_{t_{k-1}})$
$=\prod^{D}_{d=1}\tilde{p}^{d}_{t_{k}|t_{k-1}}(z^{d}_{t_{k}}|z^{1:D}_{t_{k-1}})$
$=\prod^{D}_{d=1}\delta(z^{d}_{t_{k-1}},z^{d}_{t_{k}})+R^{\theta,d}_{t}(z^{1:D}_{t_{k-1}},z^{d}_{t_{k}})\,\Delta t_{k}$ (30)
$=\underbrace{\delta(z^{1:D}_{t_{k-1}},z^{1:D}_{t_{k}})}_{=I}\;+\;\Delta t_{k}\underbrace{\sum_{d=1}\delta\left(z^{1:D\backslash d}_{t_{k-1}},z^{1:D\backslash d}_{t_{k}}\right)R^{\theta,d}_{t}(z^{1:D}_{t_{k-1}},z^{d}_{t_{k}})}_{=R^{\theta}_{t_{k-1}}}\;+\;O(\Delta t_{k}^{2}),$
where in Eq. 30 we have the approximated transition rate matrix is computed according to Eq. 13 but using $p^{\theta}_{1|t}(z^{d}_{1}|z^{1:D}_{t})$ instead of $p_{1|t}(z^{d}_{1}|z^{1:D}_{t})$.
@@
$\left\|\nu\mathcal{R}^{\theta}_{k}-\nu\mathcal{R}^{\theta,E}_{k}\right\|_{\text{TV}}=O(\Delta t_{k}^{2}).$ (31)
@@
$\|p(y^{1:D}_{K})-p_{\text{data}}\|_{\text{TV}}$ $\leq\sum^{K}_{k=1}\left(U_{k}ZD\Delta t_{k}\;+\;B_{k}(ZD\Delta t_{k})^{2}\;+\;O(\Delta t^{2}_{k})\right)$
@@
where $\Delta t=\sup_{k}\Delta t_{k}$ is the maximum step size, $\sum^{K}_{k=1}\Delta t_{k}=1$, $U=\sup_{k}U_{k}$ and $B=\sup_{k}B_{k}$.
C.2 Graph Specific Theoretical Results

### v3

@@ D.1.2 Bounded Deviation of the Generated Distribution
We start the proof by clarifying the notation for the Euler sampling approximation process. We denote its discretization timesteps by $10=t_{0}<t_{1}<\ldots<t_{K}=$ 1, with $\Delta t_{k}=t_{k}-t_{k-1}$. It is initiated at the same limit distribution as the groundtruth CTMC, $p_{\epsilon}$, and the bound to be proven will quantify the deviation that the approximated procedure incurs in comparison to the groundtruth CTMC. To accomplish so, we define $\mathcal{R}^{\theta,E}_{k}=\mathcal{R}^{\theta,E}_{t_{k-1}\to t_{k}}$ as the Markov kernel that corresponds to applying Euler sampling with the approximated rate matrix $R^{\theta}_{t}$, moving from $t_{k-1}$ to $t_{k}$. Therefore, $\mathcal{R}^{\theta,E}=\mathcal{R}^{\theta,E}_{1}\,\mathcal{R}^{\theta,E}_{2}\,\ldots\,\mathcal{R}^{\theta,E}_{K}$ and $p(y^{1:D}_{K})={\mathcal{R}^{\theta,E}}^{T}p_{\epsilon}$.
We first apply the same decomposition to the left-hand side of Theorem 8, as (Campbell et al., 2022), Theorem 1:
$\|p(y^{1:D}_{K})-p_{\text{data}}\|_{\text{TV}}$ $=\|{\mathcal{R}^{\theta,E}}^{\top}p_{\epsilon}-p_{\text{data}}\|_{\text{TV}}$
$\leq\|{\mathcal{R}^{\theta,E}}^{\top}p_{\epsilon}-\mathbb{P}_{1|0}^{\top}p_{\epsilon}\|_{\text{TV}}+\|p_{\epsilon}-\underbrace{p(z^{1:D}_{0})}_{=p_{\epsilon}}\|_{\text{TV}}$ (21)
$\leq\|{\mathcal{R}^{\theta,E}}^{\top}p_{\epsilon}-\mathbb{P}_{1|0}^{\top}p_{\epsilon}\|_{\text{TV}}$
@@
where, in Eq. 21, $\mathbb{P}_{1|0}$ denotes the path measure of the exact groundtruth CTMC and the difference between limit distributions (second term from Eq. 21) is zero since in flow matching the convergence to the limit distribution via linear interpolation is not asymptotic, in constrast to diffusion models (Austin et al., 2021), but actually attained at $t=0$. In Eq. 22, we introduce the stepwise path measure, i.e., $\mathcal{P}_{k}=\mathbb{P}_{t_{k}|t_{k-1}}$, such that $\mathbb{P}_{T|0}=\mathcal{P}_{1}\mathcal{P}_{2}\ldots\mathcal{P}_{K}$. Therefore, finding the intended upper bound amounts to establish bounds on the total variation distance for each interval $[t_{k-1},t_{k}]$.
@@
$\|{\mathcal{R}^{\theta,E}_{k}}^{\top}\nu-\mathcal{P}_{k}^{\top}\nu\|_{\text{TV}}$ $\leq\|{\mathcal{R}^{\theta,E}_{k}}^{\top}\nu-{\mathcal{R}^{\theta}_{k}}^{\top}\,\nu+{\mathcal{R}^{\theta}_{k}}^{\top}\,\nu-\mathcal{P}_{k}^{\top}\,\nu\|_{\text{TV}}$
$\leq\|\mathcal{P}_{k}^{\top}\nu-{\mathcal{R}^{\theta}_{k}}^{\top}\nu\|_{\text{TV}}+\|{\mathcal{R}^{\theta}_{k}}^{\top}\nu-{\mathcal{R}^{\theta,E}_{k}}^{\top}\nu\|_{\text{TV}},$ (23)
@@
For the first term, we use Proposition 5 from (Campbell et al., 2022) to relate the total variation distance imposed by the Markov kernels with the difference between the corresponding rate matrices:
$\|\mathcal{P}_{k}^{\top}\nu-{\mathcal{R}^{\theta}_{k}}^{\top}\nu\|_{\text{TV}}$$\leq$ $\int_{t_{k-1}}^{t_{k}}\sup_{z^{1:D}_{t}\in\mathcal{Z}^{D}}\left\{\sum_{z^{1:D}_{t+\mathrm{d}t}\neq z^{1:D}_{t}}\left|R_{t}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})-R_{t_{k-1}}^{\theta}(z^{1:D}_{t},z^{1:D}_{t+\mathrm{d}t})\right|\right\}\mathrm{d}t$
@@
where, in Eq. 24, we use the Mean Value Theorem, with $t_{c}\in(t_{k-1},t_{k})$; in Eq. 25, we use the fact that there are $ZD$ values of $z^{1:D}_{t+\mathrm{d}t}$ that differ at most in only one coordinate from $z^{1:D}_{t}$; in Eq. 26, we use the result from Proposition 7 to upper bound the time derivative of the multivariate unconditional rate matrix; and finally, in Eq. 27, we define $B_{k}=\sup_{t\in(t_{k-1},t_{k})}B_{t}$.
@@
$U^{z^{1:D}_{t}\to z^{1:D}_{t+\mathrm{d}t}}_{k}=\sqrt{C_{0}+C_{1}\mathbb{E}_{p_{1}(z^{1:D}_{1})}\left[p_{t|1}(z^{1:D}_{t}|z^{1:D}_{1})\sum^{D}_{d=1}-\log p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})\right]},$
@@
It remains to bound the second term from Eq. 23. We start by analyzing the Markov kernel $\mathcal{R}^{\theta}_{k}$ corresponding to a Markov chain with constant rate matrix $R^{\theta}_{t_{k-1}}$ between $t_{k-1}$ and $t_{k}$. In that case, from Eq. 19 we obtain:
$\mathcal{R}^{\theta}_{k}$ $=e^{R^{\theta}_{t_{k-1}}\Delta t_{k}}$
@@
$\tilde{p}_{t_{k}|t_{k-1}}$ $(z^{1:D}_{t_{k}}|z^{1:D}_{t_{k-1}})=$
$=\prod^{D}_{d=1}\tilde{p}^{(d)}_{t_{k}|t_{k-1}}(z^{(d)}_{t_{k}}|z^{1:D}_{t_{k-1}})$
$=\prod^{D}_{d=1}\delta(z^{(d)}_{t_{k-1}},z^{(d)}_{t_{k}})+R^{\theta,(d)}_{t}\left(z^{1:D}_{t_{k-1}},z^{(d)}_{t_{k}}\right)\,\Delta t_{k}$ (30)
$=\delta\left(z^{1:D}_{t_{k-1}},z^{1:D}_{t_{k}}\right)\;+\;\Delta t_{k}\sum_{d=1}\delta\left(z^{1:D\setminus(d)}_{t_{k-1}},z^{1:D\setminus(d)}_{t_{k}}\right)R^{\theta,(d)}_{t}\left(z^{1:D}_{t_{k-1}},z^{(d)}_{t_{k}}\right)\;+\;O({\Delta t_{k}}^{2}),$
where in Eq. 30 we have that the approximated transition rate matrix is computed according to Eq. 11 but using $p^{\theta,{(d)}}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$ instead of $p^{(d)}_{1|t}(z^{(d)}_{1}|z^{1:D}_{t})$. To obtain $\mathcal{R}^{\theta,E}_{k}$, we just need to convert $\tilde{p}_{t_{k}|t_{k-1}}(z^{1:D}_{t_{k}}|z^{1:D}_{t_{k-1}})$ into matrix form. Note that:
  • $\delta(z^{1:D}_{t_{k-1}},z^{1:D}_{t_{k}})$ corresponds to $I$ in matrix form;
  • From Eq. 10, $\sum_{d=1}\delta\left(z^{1:D\setminus(d)}_{t_{k-1}},z^{1:D\setminus(d)}_{t_{k}}\right)R^{\theta,(d)}_{t}\left(z^{1:D}_{t_{k-1}},z^{(d)}_{t_{k}}\right)$ corresponds to $R^{\theta}_{t_{k-1}}$ in matrix form.
These correspondences yield:
$\mathcal{R}^{\theta,E}_{k}=I+\Delta t_{k}\,R^{\theta}_{t_{k-1}}+O\left({\Delta t_{k}}^{2}\right)$
@@
$\left\|{\mathcal{R}^{\theta}_{k}}^{\top}\nu-{\mathcal{R}^{\theta,E}_{k}}^{\top}\nu\right\|_{\text{TV}}\in O\left({\Delta t_{k}}^{2}\right).$ (31)
@@
$\|p(y^{1:D}_{1})-p_{\text{data}}\|_{\text{TV}}$ $\leq\sum^{K}_{k=1}\left(U_{k}ZD\Delta t_{k}\;+\;B_{k}(ZD\Delta t_{k})^{2}\;+\;O({\Delta t_{k}}^{2})\right)$
@@
where $\Delta t=\sup_{k}\Delta t_{k}$ is the maximum step size, $\sum^{K}_{k=1}\Delta t_{k}=1$, $U=\sup_{k}U_{k}$ and $B=\sup_{k}B_{k}$. ∎
Below we provide a corollary of Theorem 8, which consists of its instantiation for graphs.
Corollary 0 (Bounded deviation of the generated graph distribution).
Let $\{G_{t}\}_{t\in[0,1]}$ be a CTMC over graphs starting with $p_{0}(G_{0})=p_{\epsilon}$ and ending with $p_{1}(G_{1})=p_{\text{data}}$, whose groundtruth rate matrix is $R_{t}$. Additionally, let $(\bar{G}_{k})_{k=0,1,\ldots,K}$ be a Euler sampling approximation of that CTMC. Then, we have:
$\|p(G_{1})-p_{\text{data}}\|_{\text{TV}}\leq\bar{U}\left(XN+E\frac{N(N-1)}{2}\right)\;+\;\bar{B}\left(XN+E\frac{N(N-1)}{2}\right)^{2}{\Delta t}+\;O(\Delta t)$
where $\bar{B}$ is defined similarly to $B$ from Theorem 8 but for graphs: $\bar{B}=\underset{\begin{subarray}{c}t\in[0,1]\end{subarray}}{\sup}\;\underset{\begin{subarray}{c}x_{1}^{(n)}\in\mathcal{X},\,e_{1}^{(ij)}\in\mathcal{E}\\ \,G_{t}\end{subarray}}{\sup}\{\bar{B}^{(n)}_{t},\bar{B}^{(ij)}_{t}\}$, with $\bar{B}^{(n)}_{t},\bar{B}^{(ij)}_{t}>0$ defined according to 6:
$p_{1|t}(x^{(n)}_{1}|G_{t})\leq\bar{B}^{(n)}_{t}\,p_{t|1}(x^{(n)}_{t}|x^{(n)}_{1})^{2}$
$p_{1|t}(e^{(ij)}_{1}|G_{t})\leq\bar{B}^{(ij)}_{t}\,p_{t|1}(e{(ij)}_{t}|e^{(ij)}_{1})^{2},$
and
$U=\underset{\begin{subarray}{c}t\in[0,1],\;\\ G_{t},\;G_{t+\mathrm{d}t}\in\mathcal{Z}^{D}\end{subarray}}{\sup}\Bigg{(}\bar{C}_{0}+\bar{C}_{1}\,\mathbb{E}_{p_{1}(G_{1})}$ $\left[p_{{t|1}}(G_{t}|G_{1})\sum^{N}_{n=1}-\log p^{\theta,{(n)}}_{1|t}(x^{(n)}_{1}|G_{t})\right]$
$+\bar{C}_{2}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{1\leq i<j\leq N}-\log p^{\theta,{(ij)}}_{1|t}(e^{(ij)}_{1}|G_{t})\right]\Bigg{)}^{\frac{1}{2}}$
Proof.
Considering the definition of a graph as $G=\left(x^{1:n:N,e^{1\leq i<j\leq N}}\right)$, we proceed through the different steps of the proof of Theorem 8.
In particular, for the discretization error, we obtain:
$\int_{t_{k-1}}^{t_{k}}\sup_{G_{t}}\left\{\sum_{G_{t+\mathrm{d}t}\neq G_{t}}\left|R_{t}(G_{t},G_{t+\mathrm{d}t})-R_{t_{k-1}}(G_{t},G_{t+\mathrm{d}t})\right|\right\}\mathrm{d}t=\bar{B}_{k}\left(XN+E\frac{N(N-1)}{2}\right)^{2}{\Delta t_{k}}^{2},$
since there are $XN+E\frac{N(N-1)}{2}$ values of $G_{t+\mathrm{d}t}$ that differ at most in only one coordinate (node or edge) from $G_{t}$, with $\bar{B}_{k}=\sup_{t\in(t_{k-1},t_{k})}\bar{B}_{t}$.
For the estimation error:
$\int_{t_{k-1}}^{t_{k}}\sup_{G_{t}}$$\left\{\sum_{G_{t+\mathrm{d}t}\neq G_{t}}\left|R_{t_{k-1}}(G_{t},G_{t+\mathrm{d}t})-R_{t_{k-1}}^{\theta}(G_{t},G_{t+\mathrm{d}t})\right|\right\}\mathrm{d}t\leq\bar{U_{k}}\left(XN+E\frac{N(N-1)}{2}\right)\Delta t_{k},$
since, again, there are $XN+E\frac{N(N-1)}{2}$ values of $G_{t+\mathrm{d}t}$ that differ at most in only one coordinate from $G_{t}$ and $U_{k}=\underset{\begin{subarray}{c}t\in[t_{k-1},t_{k}],\;\\ G_{t},\;G_{t+\mathrm{d}t}\end{subarray}}{\sup}\bar{U}^{G_{t}\to G_{t+\mathrm{d}t}}_{k}$, where:
$\bar{U}^{G_{t}\to G_{t+\mathrm{d}t}}_{k}=\Bigg{(}\bar{C}_{0}+\bar{C}_{1}\,\mathbb{E}_{p_{1}(G_{1})}$ $\left[p_{{t|1}}(G_{t}|G_{1})\sum^{N}_{n=1}-\log p^{\theta,{(n)}}_{1|t}(x^{(n)}_{1}|G_{t})\right]$
$+\bar{C}_{2}\,\mathbb{E}_{p_{1}(G_{1})}\left[p_{{t|1}}(G_{t}|G_{1})\sum_{1\leq i<j\leq N}-\log p^{\theta,{(ij)}}_{1|t}(e^{(ij)}_{1}|G_{t})\right]\Bigg{)}^{\frac{1}{2}},$
i.e., the square root of the right-hand side of Corollary .
The term $\|{\mathcal{R}^{\theta}_{k}}^{\top}\nu-{\mathcal{R}^{\theta,E}_{k}}^{\top}\nu\|_{\text{TV}}$ remains $O\left({\Delta t_{k}}^{2}\right)$.
Therefore, by finally aggregating the terms above as in Theorem 8, we obtain:
$\|p(G_{1})-p_{\text{data}}\|_{\text{TV}}\leq\bar{U}\left(XN+E\frac{N(N-1)}{2}\right)\;+\;\bar{B}\left(XN+E\frac{N(N-1)}{2}\right)^{2}{\Delta t}+\;O(\Delta t)$
D.1.3 Critical Analysis and Positioning of Corollary 1 and Corollary 2
Corollary 1 establishes that minimizing the cross-entropy (CE) loss directly corresponds to minimizing an upper bound on the rate matrix estimation error. This result provides a direct and principled justification for using the CE loss, as it promotes accurate sampling from the underlying CTMC.
Prior work derive an ELBO loss and motivate using only the CE loss based on approximations upon the derived ELBO by dropping the rate matrix-dependent terms (see Appendix C.2 in (Campbell et al., 2024)). In contrast, our bounds do not require such simplifications and are agnostic to rate matrix design choices (e.g., stochasticity level), reinforcing our training-sampling decoupling claim. Additionally, we note that concurrent work also derives a tractable ELBO (Shaul et al., 2025); however, it addresses a different family of conditional rate matrices.
Corollary 2 complements our theoretical analysis by bounding the deviation of the generated distribution from the target distribution, combining two sources of error: the discretization error from CTMC sampling, which is $\mathcal{O}(\Delta t)$, and the rate matrix estimation error, captured by Corollary 1. This type of bound is a standard objective in generative modeling, aiming to guarantee that the generation process remains faithful to the underlying data distribution. For instance, Theorem 1 in Campbell et al. (2022) presents an analogous bound in the context of discrete diffusion, though with significant differences (e.g., they assume a bounded rate matrix estimation error, while we explicitly prove it).
D.2 Theoretical Results on Architectural Expressivity